What quantity of 70 per cent acid solution must be mixed with a 30 solution to produce 480 mL of a 50 per cent solution?
Answer by mananth(16946)

70% - x ml
30% - 480-x
Mixture = 50% - 480 ml
...
0.7x+0.3(480-x)=0.5*480
0.7x+144-0.3x=240
0.4x=240-144
0.4x=96
x= 96/0.4
x=240 ml ----- 70% solution